OpenStudy (anonymous):

If 500 students attend Washington Middle School, how many are going to the mountains for vacation if the percentage going is 12%?

OpenStudy (anonymous):

60 students

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

hello

jimthompson5910 (jim_thompson5910):

12% of 500 = (12/100)*500 12% of 500 = (0.12)*500 12% of 500 = 60

OpenStudy (anonymous):

thanks for showing the work jim

jimthompson5910 (jim_thompson5910):

yw